OpenStudy (anonymous):

Given that tan(A+B)=45 deg. so find the value of (1+tanA)(1+tanB)....

OpenStudy (anonymous):

2

OpenStudy (anonymous):

Here I think your question is somewhat wrong: \[\tan(A+B) = 1\] It should be..

OpenStudy (anonymous):

or A+B=45

OpenStudy (anonymous):

Yes..

OpenStudy (anonymous):

tan(A+B)=tan45 would be the right question

OpenStudy (anonymous):

question is alright.....and the answer is 2 ...BUT HOW???? @abhiprabhu

OpenStudy (anonymous):

The question is not right..

hartnn (hartnn):

the correct question is tan(A+B)=tan45...for sure and here's the solution tan(A+B)=(tanA+tanB)/(1-tanA*tanB)=tan45=1 so,tanA+tanB=1-tanA*tanB tanA+tanB+tanA*tanB=1 1+tanA+tanB+tanA*tanB=2 (1+tanA)(1+tanB)=2